# Thermalization and Many-Body Zeno Effect in monitored Hamiltonian Dynamics

### v1

@@ Abstract
Random quantum states are essential for various applications in quantum information science. Prior approaches of generating genuine random states rely on a large bath to thermalize the system, such that a subsequent measurement on the bath post-selects a random state for the system. To reduce the size of the required bath, we propose an alternative approach based on holographic deep thermalization driven by Hamiltonian evolution, combined with mid-circuit measurements. By trading spatial and time resources, our approach achieves genuine randomness with a bath of constant size that is independent of the system size. We quantify randomness with the frame potential and analyze its dynamics throughout the evolution. Given a total evolution time, as we increase the number of mid-circuit measurements, the frame potential initially decreases exponentially with the number of measurements, due to the mechanism of holographic deep thermalization. Past a critical number of mid-circuit measurements, the frame potential rises again, signaling the onset of the quantum Zeno effect. We provide analytical results for the asymptotic behavior of the frame potential, which are in good agreement with the numerical simulations. Our findings offer practical guidance for generating Haar-random ensembles through Hamiltonian evolution and controlled measurement.
@@ I Introduction
The second law of thermodynamics states that macroscopic complex systems thermalizes–their entropy increases towards its maximum value as time evolves. To resolve its contradiction to the unitarity of quantum physics, numerous efforts have been dedicated to describe equilibration and thermalization of isolated quantum system [1, 2, 3, 4, 5], where chaotic quantum dynamics drives the isolated system to an equilibrium such that local observables appear thermal. The key is that a small system interacts with a large bath system and becomes highly entangled with the bath [6], and at the same time the reduced state of the small system appears thermal. More recently, the notion of deep thermalization [7, 8] provides a further interpretation of such a process via considering quantum measurement on the bath.
Deep thermalization concerns the state ensemble generated by quantum measurement on the bath, given that the system and bath are jointly entangled in a pure state. Conditioned on the measurement result, the system is in a random pure state; collecting conditional states under the different measurement outcomes, one therefore obtains an ensemble of pure states for the system. Refs. [7, 8] show that such measurement-induced ensemble becomes Haar random, characterized by high-order moments (hence the name ‘deep thermalization’), when the system-bath unitary dynamics are sufficiently complex. While a large bath compared to the system is essential for deep thermalization [7, 8] and quantum thermalization [1, 2, 3, 4, 5], a small bath with certain dynamical control also plays an important role in statistical physics.
The most famous example is the ‘Maxwell demon’, a small agent that measures the speed of a particle and selects high energy particles to go through to the hotter side of the system. While originally proposed as a conceptual challenge to statistical physics, it also provides a case where a small agent, with the capability of measurement and control, can change the equilibration of a large system. The essential part to resolve the seemingly violation of second law of thermodynamics is that the agent needs to be reset via dissipation to external environment. One can also provide a description of Maxwell’s demon by the quantum collision model [9], where unitary interactions and system reset alternates. Similar to Maxwell demon, stroboscopic thermalization adopts periodically driven and dissipated small ancilla to thermalize the system [10]. While the quantum collision model [9] establishes the general theory model and the corresponding dynamical equations of quantum trajectory, the generation of an ensemble of quantum states are not well understood.
Holographic deep thermalization (HDT) [11] extends the Maxwell’s demon type of bath-system interaction to the state ensemble generation. Via repeated interaction, measurement and reset, a small bath system still enables the generation of Haar random pure states, allowing a reduction of quantum memory in genuine random state generation. In the initial proposal [11], however, each unitary interaction between the system and bath is assumed to be a fixed unitary sampled from Haar ensemble. The initial theory and experimentation in Ref. [11] focuses on quantum circuit dynamics and the Hamiltonian realization of HDT is not well understood. Indeed, in a Hamiltonian dynamics, the effective realization of each Haar-typical unitary requires a long evolution time and therefore potentially creates a challenge in experiments.
Figure 1: Schematic of the quantum circuit for holographic deep thermalization with Hamiltonian dynamics and multiple mid-circuit measurements.
@@
In this work, we extend the study of holographic deep thermalization to Hamiltonian dynamics and uncover a subtle interplay between thermalization and the quantum Zeno effect [12, 13, 14, 15, 16, 17]. By employing an interacting Hamiltonian and performing intermediate measurements on the bath, we aim to generate Haar-random states. Our findings show that increasing the number of intermediate measurements generally enhances the randomness of the resulting state ensemble. However, beyond a certain threshold, excessive measurements trigger the quantum Zeno effect, inhibiting the system’s evolution and diminishing the ensemble’s randomness. We provide both analytical and numerical insights to determine the optimal number of measurements that maximizes Haar-randomness, given a fixed total evolution time.
@@ II Overview
Hamiltonian evolution can also be used to generate Haar-random states and is often more practical to implement in quantum simulation platforms. In our approach, we utilize a Hamiltonian that satisfies the eigenstate-thermalization-hypothesis (ETH) to drive the system toward Haar randomness. The entire system consists of the system with $n_{\rm s}$ qubits and the bath (ancillary) with $n_{\rm b}$ qubits. The total evolution time is fixed at $T$, during which we insert $n$ intermediate measurements on the bath qubits. As a result, each segment of unitary evolution lasts for $\Delta t=T/n$. The bath is measured and then reset after each segment. A schematic of the quantum circuit is shown in Fig. 1. Different measurement outcomes yield different output states, collectively forming an ensemble. We adopt the frame potential to characterize the randomness of the ensemble quantitatively.
For the Hamiltonian evolution process shown in Fig. 1, the frame potential initially decreases and then increases depending on the number of measurements $n$, as illustrated by the blue curves in Fig. 2. The initial decrease in the small-$n$ regime is well captured by the HDT theory, shown as the orange curves in Fig. 2. In contrast, the increase in the large-$n$ regime can be explained by the quantum Zeno effect, represented by the green curves. These theoretical descriptions hold across different bath sizes, as demonstrated by the comparison between cases with a small number of bath qubits in Figs. 2(a)(d) and those with a larger number of bath qubits in Figs. 2(b)(c).
When the total evolution time $T$ is large, the frame-potential decrease saturates once $n$ approaches $n_{\rm sat}\sim O\left(\left(n_{\rm s}-{\rm log}_{2}r\right)/n_{\rm b}\right)$, as indicated by the orange dashed lines in Fig. 2(a)(b), where $r$ is a Hamiltonian dependent parameter. Beyond this point, the frame potential first saturates to a plateau, with fluctuations induced by dynamical revival from finite-system-size effect (see Appendix A). After the plateau, the frame potential begins to increase due to the onset of the quantum Zeno effect, occurring around $n_{\rm Zeno}\sim O\left(\sqrt[\alpha]{T^{\alpha+1}/n_{\rm s}}\right)$, as indicated by the green dashed lines in Fig. 2(a)(b), where $\alpha$ is another Hamiltonian depended parameter.
For small $T$ with $n_{\rm sat}>n_{\rm Zeno}$ (Fig. 2(c)(f)), HDT and the Zeno effect compete, and the minimum frame potential is reached near $n_{\gamma}\sim O\left(T/\sqrt[\alpha+1]{n_{\rm b}}\right)$ which reflects the optimal balance between thermalization and measurement-induced suppression.
@@ III Set-up
In HDT, both the system and bath are initialized in a trivial state that is easy to prepare; here, we choose the all-zero state, denoted as $|\psi_{0}\rangle\otimes|\phi_{0}\rangle$. The system and bath are then entangled through evolution under a time-independent Hamiltonian $H$. They jointly evolve for a time interval $\Delta t$ according to the Schrödinger equation. For brevity, we set $\hbar=1$. After each evolution step, the bath is measured in the computational basis $|\phi_{m}\rangle$} and reset to the trivial state $|\phi_{0}\rangle$. The cycle of Hamiltonian evolution, measurement, and reset is repeated $n$ times, with the total evolution time fixed at $T$. At the end of the process, the system collapses into one of the possible states $|\psi_{z}\rangle$ with probability $p_{z}$, where $z$ indexes all possible measurement trajectories. These outcomes define an ensemble $\mathcal{E}=\{p_{z},|\psi_{z}\rangle\}$.
Figure 2: Frame potential as a function of the number of measurements. The blue lines represent numerical results, while the orange and green lines correspond to analytical results for the asymptotic region of HDT in Eq. (7) and the Zeno effect in Eq. (12), respectively. Other parameters are: (a),(d) 5 system qubits and 3 bath qubits; (b),(c),(e),(f) 7 system qubits and 1 bath qubit. Total evolution time is $T=15$ for (a),(b),(d),(e) and $T=5$ for (c),(f). The orange, green, and blue dashed lines indicate the positions of $n_{\rm sat}$, $n_{\rm Zeno}$, and $n_{\gamma}$, respectively, with $r=0.1$
The randomness of the ensemble $\mathcal{E}$ can be quantified using the frame potential, defined as
$F^{(K)}=\sum_{z,z^{\prime}}p_{z}p_{z^{\prime}}\left|\langle\psi_{z}|\psi_{z^{\prime}}\rangle\right|^{2K}.$ (1)
In the ideal case, the Haar random frame potential is
$F_{\rm Haar}^{(K)}=\frac{(N_{\rm s}-1)!K!}{(N_{\rm s}+K-1)!},$ (2)
@@
In prior work on DT, the transverse-field Ising model has been used to approximate Haar-random dynamics. In our approach, we adopt the same Hamiltonian to implement HDT, which is given by
$H_{\rm Ising}=J_{x}\sum\sigma_{j}^{x}+J_{z}\sum\sigma_{j}^{z}+J_{zz}\sum\sigma^{z}_{j}\sigma_{j+1}^{z},$ (3)
@@
$H_{YY}$ $=$ $H_{\rm Ising}+J_{yy}\sum\sigma^{y}_{j}\sigma_{j+1}^{y},$ (4)
$H_{XXX}$ $=$ $H_{\rm Ising}+J_{xxx}\sum\sigma_{j-1}^{x}\sigma^{x}_{j}\sigma_{j+1}^{x}.$ (5)
In the context of Hamiltonian evolution, we consider a system of qubits where a subset of them is treated as a bath and measured, while the remaining qubits form an ensemble that approaches Haar-random. Generally, the Hamiltonian can be written into three parts
$H$ $=H_{{\rm s}}+H_{{\rm b}}+H_{{\rm c}},$ (6)
where $H_{{\rm s}},H_{{\rm b}}$ and $H_{{\rm c}}$ correspond to the system, bath and interaction Hamiltonian.
@@ IV Holographic deep thermalization
Figure 3: The convergence dynamics of frame potential as a function of the number of measurements for (a) $H_{\rm Ising}$, (b) $H_{YY}$, and (c) $H_{XXX}$. y axis is the frame potential difference with the converged value in Eq. (8). Blue lines represent numerical results, while orange lines correspond to analytical predictions for HDT. The total evolution time is $T=15$. The system consists of 7 qubits, with 1 bath qubit.
In the context of holographic deep thermalization, the insertion of multiple mid-circuit measurements significantly accelerates the thermalization process, potentially leading to exponential speedup in approaching Haar randomness. If the unitaries are randomly sampled from the Haar measure, the expected first-order frame potential in holographic deep thermalization is given by
@@
where the corrected saturation value is
$q_{\rm 1}=\frac{N_{\rm s}^{2}(N_{\rm b}+1)}{N_{\rm s}^{2}N_{\rm b}+1}F_{\rm Haar}^{(1)}.$ (8)
For the higher order frame potential, a lower bound can be obtained as
$F^{(K)}=(1-q_{K})\left(\frac{N_{\rm s}^{2}N_{\rm b}}{N_{\rm s}^{2}N_{\rm b}^{2}-1}\right)^{n}-q_{K},$ (9)
where the factor ($K\geqslant 2$) is
$q_{K}=\left(1+\frac{2^{K}-1}{N_{\rm b}}\right)F_{\rm Haar}^{(K)}.$ (10)
Although Hamiltonian evolution, which is spatially local, differs from random Haar-sampled unitaries, it agrees well with the Haar unitary results when the number of measurements $n$ is small. As illustrated by the blue dotted lines in Fig. 3, the frame potential exhibits a rapid decay in this regime. This behavior is well captured by the HDT theory described in Eq. (7), which predicts an exponential decrease, as shown by the orange lines in Fig. 3. Notably, this agreement holds across different Hamiltonians, as demonstrated in Fig. 3(a)(b)(c).
Phenomenologically, when the frame potential approaches the same order as its minimal value, $F^{(1)}=(r+1)F_{\rm c}$, saturation starts to kick in. The number of measurements required for saturation is
$n_{\rm sat}\approx\frac{n_{\rm s}-{\rm log}_{2}r}{n_{\rm b}}\leavevmode\nobreak\ .$ (11)
@@
Figure 4: Frame potential for $K$-designs corresponding to (a) $K=1$, (b) $K=3$, and (c) $K=10$. The black solid line represents the analytical approach. Blue dots indicate results from the non-integrable Hamiltonian ($J_{z}\neq 0$), while green dots correspond to the integrable Hamiltonian ($J_{z}=0$).
@@
We find that both integrable and non-integrable local Hamiltonians remain far from the minimum value predicted by HDT theory, as shown in Fig. 4(c). This deviation is likely due to the inherent locality of the Hamiltonian. Although a chaotic Hamiltonian may exhibit ergodic behavior in local observables, it does not fully explore the entire Hilbert space. One fundamental limitation is energy conservation, which restricts dynamics to a finite energy shell. Moreover, due to locality constraints, the evolution cannot uniformly cover Hilbert space in the same way as a Haar random unitary. Furthermore, the Hamiltonian contains only a limited number of independent parameters, meaning that the matrix elements of the evolution operator $U=e^{-iH\Delta t}$ are not fully independent. In contrast, a Haar-random unitary has a number of independent parameters equal to the dimension of the Hilbert space minus one. Consequently, the dynamics generated by a local Hamiltonian cannot reach the level of randomness characteristic of a typical Haar-random unitary. These observations highlight the intrinsic limitations of using local Hamiltonians and restricted resources to generate high-order quantum randomness.

### v2

# Many-Body Anti-Zeno Thermalization and Zeno Determinism in Monitored Hamiltonian Dynamics

@@ Abstract
Random quantum states are essential for quantum information science, with applications ranging from quantum computing to cryptography. Prior approaches for generating these states often rely on using a large bath to thermalize a smaller system, with a subsequent measurement on the bath used to post-select a random state. To reduce the required size of the bath, we propose a resource-efficient scheme using holographic deep thermalization driven by Hamiltonian evolution, combined with mid-circuit measurements. This scheme relies on dynamical circuits, enabling a trade-off between spatial and temporal resources and allowing the generation of genuinely random states with only a constant-size bath. We quantify the randomness using the frame potential and derive its asymptotic behavior, which shows good agreement with our numerical simulations and experimental results on IBM quantum devices. For a fixed total evolution time, increasing the number of mid-circuit measurements initially produces an exponential decrease in the frame potential—a quantum anti-Zeno behavior arising from holographic deep thermalization. Past a critical number of mid‑circuit measurements, the frame potential rises again, signaling the onset of the quantum Zeno effect.
@@ I Introduction
Quantum systems equilibrate and thermalize [reimann2008foundation, rigol2008thermalization, short2011equilibration, rigol2012alternatives, short2012quantum] through chaotic and entangling unitary interaction with a large bath [popescu2006entanglement], causing its reduced state to appear thermal [Donovan2025, Mori2023]. More recently, the notion of deep thermalization [ho2022exact, cotler2023emergent] further considers the state ensemble generated by quantum measurement on the bath, given that the system-bath state is pure. Conditioned on the measurement result, the system is in a random pure state [Giulia2025]; collecting conditional states under the different measurement outcomes, one therefore obtains an ensemble of pure states for the system. Refs. [ho2022exact, cotler2023emergent] show that a measurement-induced ensemble becomes Haar random when the unitary dynamics of the system and the bath are sufficiently intricate.
While a large bath compared to the system is essential for deep thermalization [ho2022exact, cotler2023emergent] and quantum thermalization [reimann2008foundation, rigol2008thermalization, short2011equilibration, rigol2012alternatives, short2012quantum], a small bath with certain dynamical control also plays an important role in statistical physics [metcalf2020engineered], as exemplified by the famous “Maxwell’s demon” [ciccarello2022quantum]. Via repeated interaction, measurement and reset, holographic deep thermalization(HDT) [zhang2025holographic] further allows a small bath to enable Haar random pure states on the system. In the initial proposal [zhang2025holographic], however, each unitary interaction between the system and bath is assumed to be a fixed unitary sampled from Haar ensemble. While effective for circuit-based platforms, this assumption leaves open the question of how HDT can be faithfully realized using continuous Hamiltonian dynamics. In particular, implementing Haar-typical unitaries typically requires long evolution times, which may present constraints for experimental realizations.
@@
In this work, we extend the study of holographic deep thermalization to Hamiltonian dynamics and uncover a subtle interplay between anti-Zeno thermalization and the quantum Zeno effect. By employing an interacting Hamiltonian and performing intermediate measurements on the bath, we aim to generate Haar-random states. Given a fixed evolution time, our findings show that increasing the number of intermediate measurements initially enhances the randomness of the resulting state ensemble, in analogy to the anti-Zeno effect(AZE) [kofman2000, Facchi2001, Fischer2001, Chaudhry2014, Leppenen2022, Gumber2023, Zhang2025]. However, beyond a certain threshold, excessive measurements trigger the quantum Zeno effect(ZE) [misra1977zeno, FACCHI200012, facchi2022quantum, mccusker2013experimental, schafer2014experimental, signoles2014confined, slichter2016quantum, Becker2021, zeni2025, Li2018, Biella2021], inhibiting the bath’s evolution and diminishing the ensemble’s randomness. We provide both analytical and numerical insights to determine the optimal number of measurements that maximizes Haar-randomness for a fixed evolution duration. Our results generalize AZE and ZE to partial mid-circuit measurement and connect them to deep thermalization.
Figure 1: Schematic of the quantum circuit implementing holographic deep thermalization via Hamiltonian dynamics with multiple mid-circuit measurements.
@@ II Overview
We utilize a Hamiltonian that satisfies the eigenstate thermalization hypothesis (ETH) to drive the system toward Haar randomness. The full system comprises a system register of $n_{\rm s}$ qubits and a bath of $n_{\rm b}$ ancillary qubits, for a total of $n_{\rm all}=n_{\rm s}+n_{\rm b}$ qubits. As shown in Fig. 1, the total evolution time is fixed at $T$, during which we insert $n$ intermediate measurements on the bath qubits. The bath is measured and then reset after each segment. We quantify the randomness of the measurement-generated pure-state ensemble using the frame potential.
For the Hamiltonian evolution process shown in Fig. 1, the frame potential initially decreases and then increases depending on the number of measurements $n$, as illustrated by the blue curves in Fig. 2. The initial decrease in the small-$n$ AZE regime is well captured by the HDT theory (orange curves). In contrast, the increase in the large-$n$ regime can be explained by the quantum ZE (green curves). These theoretical descriptions hold across different bath sizes, as demonstrated by the comparison between cases with a small number of bath qubits in Figs. 2(a)(d) and those with a larger number of bath qubits in Figs. 2(b)(c).
When the total evolution time $T$ is large, the frame‑potential decrease saturates once $n$ approaches $n_{\rm sat}\sim O\left(\left(n_{\rm s}-{\rm log}_{2}r\right)/n_{\rm b}\right)$, as indicated by the orange dashed lines in Fig. 2(a)(b), where $r$ is a Hamiltonian dependent parameter. Beyond this point, the frame potential first saturates to a plateau, with fluctuations induced by dynamical revival from finite-system-size effect (Appendix A). After the plateau, the frame potential begins to increase due to the onset of the quantum ZE, occurring around $n_{\rm Zeno}\sim O\left(\sqrt[\alpha]{T^{\alpha+1}/n_{\rm s}}\right)$, as indicated by the green dashed lines in Fig. 2(a)(b), where $\alpha$ is another Hamiltonian-dependent parameter.
For small $T$ with $n_{\rm sat}>n_{\rm Zeno}$ (Fig. 2(c)(f)), HDT and the ZE compete, and the minimum frame potential is reached near $n_{\gamma}\sim O\left(T/\sqrt[\alpha+1]{n_{\rm b}}\right)$ which reflects the optimal balance between thermalization and measurement-induced suppression.
Figure 2: Frame potential as a function of the number of measurements. The blue lines represent numerical results, while the orange and green lines correspond to analytical results for the asymptotic region of HDT in Eq. (6) and the ZE in Eq. (9), respectively. Other parameters are: (a),(d) $n_{\rm s}=5$, $n_{\rm b}=3$; (b),(c),(e),(f) $n_{\rm s}=7$, $n_{\rm b}=1$. Total evolution time is $T=15/J_{zz}$ for (a),(b),(d),(e) and $T=5/J_{zz}$ for (c),(f). The orange, green, and blue dashed lines indicate the positions of $n_{\rm sat}$, $n_{\rm Zeno}$, and $n_{\gamma}$, respectively, with $r=0.1$. Other parameters are $J_{z}=0.9J_{zz}$ and $J_{x}=1.4J_{zz}$.
@@ III Set-up
In HDT, both the system and bath are initialized in a trivial state that is easy to prepare; here, we choose the all-zero state, denoted as $|\psi_{0}\rangle\otimes|\phi_{0}\rangle$. The system and bath are then entangled through evolution $U=e^{-iH\Delta t}$ under a time-independent Hamiltonian $H$ for a time interval $\Delta t$. For brevity, we set $\hbar=1$. After each evolution step, the bath is measured in the computational basis $|\phi_{m}\rangle$} and reset to the trivial state $|\phi_{0}\rangle$. The cycle of Hamiltonian evolution, measurement, and reset is repeated $n$ times, with the total evolution time fixed at $T$. At the end of the process, the system collapses into one of the possible states $|\psi_{z}\rangle$ with probability $p_{z}$, where $z$ indexes all possible measurement trajectories. These outcomes define an ensemble $\mathcal{E}=\{p_{z},|\psi_{z}\rangle\}$.
The randomness of the ensemble $\mathcal{E}$ is characterized by the $K$-th frame potential [ippoliti2023dynamical]
$F^{(K)}=\sum_{z,z^{\prime}}p_{z}p_{z^{\prime}}\left|\langle\psi_{z}|\psi_{z^{\prime}}\rangle\right|^{2K}\penalty 10000\ ,$ (1)
which measures the average overlap between different output states in the ensemble. A smaller frame potential indicates greater randomness. For comparison, the Haar ensemble provides the benchmark for maximal randomness. In this ideal case, the $K$-th frame potential takes the closed form
$F_{\rm Haar}^{(K)}=\frac{(N_{\rm s}-1)!K!}{(N_{\rm s}+K-1)!}\penalty 10000\ ,$ (2)
@@
In prior work on HDT, the transverse and longitudinal-field Ising model has been used to approximate Haar-random dynamics [cotler2023emergent]. In our approach, we adopt the same spin-chain Hamiltonian with open boundary condition to implement HDT, which is given by
$H_{\rm Ising}=J_{x}\sum_{j=0}^{n_{\rm all}-1}\sigma_{j}^{x}+J_{z}\sum_{j=0}^{n_{\rm all}-1}\sigma_{j}^{z}+J_{zz}\sum_{j=0}^{n_{\rm all}-2}\sigma^{z}_{j}\sigma_{j+1}^{z}\penalty 10000\ ,$
@@
$H_{YY}$ $=$ $H_{\rm Ising}+J_{yy}\sum_{j=0}^{n_{\rm all}-2}\sigma^{y}_{j}\sigma_{j+1}^{y}\penalty 10000\ ,$ (4)
$H_{XXX}$ $=$ $H_{\rm Ising}+J_{xxx}\sum_{j=1}^{n_{\rm all}-2}\sigma_{j-1}^{x}\sigma^{x}_{j}\sigma_{j+1}^{x}\penalty 10000\ .$ (5)
@@
Figure 3: The convergence dynamics of frame potential as a function of the number of measurements for (a) $H_{\rm Ising}$, (b) $H_{YY}$, and (c) $H_{XXX}$. Blue lines represent numerical results, while orange lines correspond to analytical predictions for HDT. Other parameters are $n_{\rm s}=7$, $n_{\rm b}=1$, $J_{z}=0.9J_{zz}$, $J_{x}=1.3J_{zz}$, and $T=15/J_{zz}$.
@@ IV Holographic deep thermalization
In the context of HDT, the insertion of multiple mid-circuit measurements significantly accelerates the thermalization process, potentially leading to exponential speedup in approaching Haar randomness. If the unitaries are randomly sampled from the Haar measure, the expected first-order frame potential in HDT is given by [zhang2025holographic]
@@
where the corrected saturation value is $q_{\rm 1}\equiv\frac{N_{\rm s}^{2}(N_{\rm b}+1)}{N_{\rm s}^{2}N_{\rm b}+1}F_{\rm Haar}^{(1)}.$ For the higher order frame potential ($K\geqslant 2$), a lower bound can be obtained as
$F^{(K)}=(1-q_{K})\left(\frac{N_{\rm s}^{2}N_{\rm b}}{N_{\rm s}^{2}N_{\rm b}^{2}-1}\right)^{n}+q_{K}\penalty 10000\ ,$ (7)
where the corrected saturation value is $q_{K}=\left(1+{(2^{K}-1)}/{N_{\rm b}}\right)F_{\rm Haar}^{(K)}.$
Although the spatially-local Hamiltonian evolution differs from random Haar-sampled unitaries, the resulting frame potential dynamics agrees well with the Haar unitary results when the number of measurements $n$ is small. As illustrated by the blue dotted lines in Fig. 3, the frame potential exhibits a rapid decay in this regime. This behavior is well captured by the HDT theory described in Eq. (6), which predicts an exponential decrease, as shown by the orange lines in Fig. 3. Notably, this agreement holds across different Hamiltonians, as demonstrated in the three subplots of Fig. 3.
Phenomenologically, when the frame potential approaches the same order as its minimal value, $F^{(1)}=(r+1)F_{\rm c}$ with small constant $r$, saturation starts to kick in. The number of measurements required for saturation is
$n_{\rm sat}\approx\frac{n_{\rm s}-{\rm log}_{2}r}{n_{\rm b}}\penalty 10000\ .$ (8)
@@
Figure 4: Frame potential for $K$-designs corresponding to (a) $K=1$, (b) $K=3$, and (c) $K=10$. The black solid line represents the analytical approach. Blue dots indicate results from the non-integrable Hamiltonian ($J_{z}=0.9J_{zz}$), while green dots correspond to the integrable Hamiltonian ($J_{z}=0$). Other parameters are $n_{\rm s}=7$, $n_{\rm b}=1$, $J_{x}=1.3J_{zz}$, and $T=15/J_{zz}$ .
@@
We find that both integrable and non-integrable local Hamiltonians remain far from the minimum value predicted by HDT theory, as shown in Fig. 4(c). This deviation is likely due to the inherent locality of the Hamiltonian. Although a chaotic Hamiltonian may exhibit ergodic behavior in local observables, it does not fully explore the entire Hilbert space. One fundamental limitation is energy conservation, which restricts dynamics to a finite energy shell. Moreover, due to locality constraints, the evolution cannot uniformly cover the Hilbert space in the same way as a Haar random unitary. Furthermore, the Hamiltonian contains only a limited number of independent parameters, meaning that the matrix elements of the evolution operator $U=e^{-iH\Delta t}$ are not fully independent. In contrast, a Haar-random unitary has a number of independent parameters equal to the dimension of the Hilbert space minus one. Consequently, the dynamics generated by a local Hamiltonian cannot reach the level of randomness characteristic of a typical Haar-random unitary. These observations highlight the intrinsic limitations of using local Hamiltonians and restricted resources to generate high-order quantum randomness.
Figure 5: Frame potential as a function of the number of measurements for (a) $H_{\rm Ising}$, (b) $H_{YY}$, and (c) $H_{XXX}$. Blue lines represent numerical results, while green lines correspond to analytical predictions for ZE with (a) $\alpha=3$, and (b)(c) $\alpha=1$. The total evolution time is $T=15/J_{zz}$. The system consists of 5 qubits, with 3 bath qubit. Other parameters are $J_{z}=0.9J_{zz}$ and $J_{x}=1.3J_{zz}$.